A Bibliography of Publications in SIGCSE Bulletin: 1960–1969

Nelson H. F. Beebe University of Utah Department of Mathematics, 110 LCB 155 S 1400 E RM 233 Salt Lake City, UT 84112-0090 USA

Tel: +1 801 581 5254 FAX: +1 801 581 4148

       E-mail: beebe@math.utah.edu, beebe@acm.org,
                  beebe@computer.org (Internet)
      WWW URL: http://www.math.utah.edu/~beebe/

21 April 2018 Version 1.02

Title word cross-reference

Arts [Ral69].

Bachelor [Ral69]. basic [Mat69]. beginning [CO69]. Buffalo [Ral69].

Carnegie [Wul69]. Carnegie-Mellon [Wul69]. Computer [Ham69, NT69, Tho69, Den69, For69, Har69, Mat69]. course [Mat69]. CSE [Kor69].

degree [Ham69, Ral69]. dirty [Har69].

education [Ham69, NT69]. electrical [Tho69]. engineering [Tho69]. examinations [For69]. Exams [Wul69]. exercise [CO69].

given [Mat69].

REFERENCES 2

higher [Ham69]. introduction [Har69]. leading [Ral69]. list [CO69]. manipulation [CO69]. masses [Har69]. mathematics [NT69]. Mellon [Wul69]. option [Tho69]. organization [Den69]. Ph.D. [For69]. practice [Kor69]. Principles [Den69]. Program [Ral69]. programs [Ham69]. project [Mat69]. qualifying [For69]. Quick [Har69]. related [Ham69]. school [NT69]. science [For69, Ham69, Mat69, Tho69]. secondary [NT69]. Stanford [For69]. students [CO69, Har69]. SUNY [Ral69]. Syllabi [For69]. system [Den69]. teachers [NT69]. term [Mat69]. theory [Kor69]. U.S [Ham69]. University [Wul69, For69, Mat69, Tho69]. Washington [Mat69]. Windsor [Tho69].

References

Channen:1969:ELM

[CO69] E. W. Channen and E. I. Organick. An exercise in list manipulation for beginning students. SIGCSE Bulletin (ACM Special Interest Group on Computer Science Education), 1(2):13–32, June 1969. CODEN SIGSD3. ISSN 0097-8418 (print), 2331-3927 (electronic).

Denning:1969:PCS

[Den69] Peter J. Denning. Principles of computer system organization. SIGCSE Bulletin (ACM Special Interest Group on Computer Science Education), 1(3):27–30, October 1969. CODEN SIGSD3. ISSN 0097-8418 (print), 2331-3927 (electronic).

Forsythe:1969:SQE

[For69] G. Forsythe. Syllabi and qualifying examinations for the Ph.D. in computer science at Stanford University. SIGCSE Bulletin (ACM Special Interest Group on Computer Science Education), 1(4):25–34, December 1969. CODEN SIGSD3. ISSN 0097-8418 (print), 2331-3927 (electronic).

Hamblen:1969:CSR

[Ham69] John W. Hamblen. Computer science and related degree programs in U.S. higher education. SIGCSE Bulletin (ACM Special Interest Group on Computer Science Education), 1(4):9–13, December 1969. CODEN SIGSD3. ISSN 0097-8418 (print), 2331-3927 (electronic).

Hart:1969:QDI

[Har69] Robert R. Hart. Quick and dirty introduction to the computer for masses of students. SIGCSE Bulletin (ACM Special Interest Group on Computer Science Education), 1(2):35, June 1969. CODEN SIGSD3. ISSN 0097-8418 (print), 2331-3927 (electronic).

Korfhage:1969:CTP

[Kor69] Robert R. Korfhage. CSE: theory or practice? SIGCSE Bulletin (ACM Special Interest Group on Computer Science Education), 1(4): 6–7, December 1969. CODEN SIGSD3. ISSN 0097-8418 (print), 2331-3927 (electronic).

Matula:1969:TPG

[Mat69] D. Matula. A term project given in the basic computer science course at Washington University. SIGCSE Bulletin (ACM Special Interest Group on Computer Science Education), 1(4):16–19, December 1969. CODEN SIGSD3. ISSN 0097-8418 (print), 2331-3927 (electronic).

Nievergelt:1969:CES

[NT69] J. Nievergelt and K. J. Travers. Computer education for secondary school mathematics teachers. SIGCSE Bulletin (ACM Special Interest Group on Computer Science Education), 1(4):7–9, December 1969. CODEN SIGSD3. ISSN 0097-8418 (print), 2331-3927 (electronic).

Ralston:1969:PLD

[Ral69] A. Ralston. Program leading to the degree of Bachelor of Arts at SUNY at Buffalo. SIGCSE Bulletin (ACM Special Interest Group on Computer Science Education), 1(4):13–14, December 1969. CODEN SIGSD3. ISSN 0097-8418 (print), 2331-3927 (electronic).

Thomas:1969:CSO

[Tho69] P. A. V. Thomas. Computer science option in electrical engineering at the University of Windsor. SIGCSE Bulletin (ACM Special Interest Group on Computer Science Education), 1(4):14–16, December 1969. CODEN SIGSD3. ISSN 0097-8418 (print), 2331-3927 (electronic).

Wulf:1969:ECM

[Wul69] W. Wulf. Exams from carnegie-mellon university. SIGCSE Bulletin (ACM Special Interest Group on Computer Science Education), 1(4): 19–25, December 1969. CODEN SIGSD3. ISSN 0097-8418 (print), 2331-3927 (electronic).